Two point charges Q1 = +5.00 nC and Q2 = -3.00 nC are separated by 35.0 cm (a) What is the electric potential at a point midway between the charges? (b) What is the potential energy of the pair of charges?
1
Expert's answer
2020-10-22T18:08:49-0400


Let "P" represent the point midway between the charges.

Let "r_1" be the distance between "Q_1" and "P" .

Let "r_2" be the distance between "Q_2" and "P" .

Let "V_p" represent the potential at "P" .

Let "k_e" represent coulomb's constant.

Let "U" represent the potential energy of the pair of charges.

"r_1=r_2=\frac{35cm}{2}=0.175m"

"\begin{aligned}\textsf{(a)}
&V_p=k_e(\frac{Q_1}{r_1}+\frac{Q_2}{r_2})\\
&V_p=(8.99×10^9N.m²/C²)×(\frac{5.00×10^{-9}C}{0.175m}-\frac{3.00×10^{-9}C}{0.175m})\\
&V_p=102.74V
\end{aligned}"


"\begin{aligned}\textsf{(b)}&U=k_e\frac{Q_1Q_2}{r_1+r_2}\\
&U=(8.99×10^9N.m²/C²)(\frac{(5.00×10^{-9})C×(-3.00×10^{-9})C}{0.175m+0.175m})\\
&U=-3.85×10^{-7}J\end{aligned}"